### v2

# Flat-band FFLO State from Quantum Geometric Discrepancy

Abstract
We propose a new scheme for the Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) Cooper pairing states within ultraflat bands. Central to our approach is the concept of “quantum geometric discrepancy” (QGD), which characterizes the discrepancy in the quantum geometry of paired electrons, giving rise to the flat-band FFLO instability. Remarkably, we find that this instability is directly related to a new quantum geometric quantity we term “anomalous quantum distance”, which formally captures QGD. To model both QGD and the anomalous quantum distance, we examine a flat-band electronic Hamiltonian with tunable spin-dependent quantum metrics. Utilizing the band-projection method, we analyze the QGD-induced FFLO instability from pair susceptibility near the superconducting critical temperature. Furthermore, we perform self-consistent mean-field calculations to obtain the phase diagram of the BCS-FFLO transition driven by QGD, which aligns well with our analytical results. We emphasize that QGD serves as a distinctive protocol for stabilizing the flat-band FFLO phase.
Introduction.—Understanding unconventional superconductivity is a central theme in modern condensed matter physics, with important implications for both theoretical and experimental aspects Sigrist and Ueda (1991); Mineev et al. (1999); Bauer and Sigrist (2012); Scalapino (2012). Among the various mechanisms that extend beyond the conventional Bardeen-Cooper-Schrieffer (BCS) theory of pairing Bardeen et al. (1957), the Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) phase is notable for its unique characterization of Cooper pairs with finite center-of-mass momentum $\bm{Q}$ Fulde and Ferrell (1964); Larkin and Ovchinnikov (1965); Kinnunen et al. (2018); Agterberg et al. (2020). Various mechanisms have been explored to realize these exotic FFLO states in systems characterized by dispersive bands Lu et al. (2015); Xi et al. (2016); Zhou et al. (2016); Saito et al. (2016); Liu (2017); Song and Koshelev (2019); Xie and Law (2023); Yuan (2023); Nakamura et al. (2024); Wan et al. (2023).
In the original formulation of the FFLO phase, a spin-singlet pairing state with finite center-of-mass momentum $\bm{Q}$ ($|\bm{Q}|\propto B/v_{F}$) is favored in the presence of a large Zeeman field $B$ [see Fig. 1(a)]. However, the applicability of this microscopic mechanism becomes problematic in flat-band systems, where the Fermi velocity $v_{F}$ is nearly zero. This results in an unphysical divergence of $\bm{Q}$ as $v_{F}\rightarrow 0$, in the presence of a finite magnetic field. Consequently, two critical questions arise: First, what mechanism can stabilize the FFLO state as $v_{F}\rightarrow 0$? Second, what determines the center-of-mass momentum $\bm{Q}$ of Cooper pairs in the flat-band FFLO state? In dispersive bands, this is typically governed by Fermi surface nesting; however, in flat bands, the poorly defined Fermi surface weakens this relationship.
Figure 1: Schematic illustrations of (a) conventional FFLO from Zeeman effect, and (b) flat-band FFLO from QGD. In (a), an external $B$ field induces spin-split Fermi surfaces, leading to the formation of Cooper pairs with finite momentum $\bm{Q}$, where $|\bm{Q}|\propto B/v_{F}$. In (b), the finite momentum $\bm{Q}$ is stabilized when the quantum metrics of the paired electrons are different ($\mathcal{G}_{\uparrow}\neq\mathcal{G}_{\downarrow}$) within flat bands.
Lying in the context of the recent studies of the pronounced quantum geometric effects in flat-band systems Peotta and Törmä (2015); Julku et al. (2016); Liang et al. (2017); Hu et al. (2019); Julku et al. (2020); Xie et al. (2020); Peri et al. (2021); He et al. (2021); Herzog-Arbeitman et al. (2022); Törmä et al. (2022, 2018); Huhtinen et al. (2022); Hofmann et al. (2023); Jiang and Barlas (2023); Chen and Huang (2023); Han et al. (2024); Ticea et al. (2024); Wang and Huang (2024); Chen and Law (2024); Hu et al. (2023); Kitamura et al. (2022); Törmä (2023); Tian et al. (2023); Yu et al. (2024), in this work, we address the aforementioned questions by linking the FFLO instability to the quantum geometry of Bloch wave functions. For concreteness, we propose a new scheme for FFLO pairing states within an isolated flat band, driven by “quantum geometric discrepancy (QGD)”. This concept refers to the discrepancy in the quantum geometry of the wave functions of the two electrons forming a Cooper pair, which explicitly breaks time reversal ($\mathcal{T}$) symmetry. Based on the band-projection formalism, we analyse the static pair susceptibility and establish that the critical temperature for finite-momentum pairing, which leads to FFLO instability, is directly associated with the band average of a quantum geometric quantity [as described in Eq. (6)]—“anomalous quantum distance (AQD)” $\bar{d}_{\bm{k}^{\prime},\bm{k}}$, formally defined as
$\bar{d}^{2}_{\bm{k}^{\prime},\bm{k}}=1-\left|\langle u_{\bm{k}^{\prime}\uparrow}|u^{*}_{-\bm{k}\downarrow}\rangle\right|^{2}.$ (1)
Here, $u_{\bm{k}\sigma}$ represents the Bloch wave function at momentum $\bm{k}$ and spin $\sigma$. AQD quantifies the overlap between the wave functions of paired electrons. In the presence of $\mathcal{T}$ symmetry, where $u_{\bm{k}\uparrow}=u^{*}_{-\bm{k}\downarrow}$, the quantity $\bar{d}_{\bm{k},\bm{k}^{\prime}}$ becomes the well-known Hilbert–Schmidt quantum distance Shapere and Wilczek (1989); Bužek and Hillery (1996); Dodonov et al. (2000); Rhim et al. (2020), which measures the quantum mechanical distance between two Bloch states. In this scenario, the BCS (zero-momentum) pairing state is favored. In the absence of $\mathcal{T}$ symmetry, $u_{\bm{k}\uparrow}\neq u^{*}_{-\bm{k}\downarrow}$ generally makes this relation unestablished, and Eq. (1) will exhibit anomalous behaviors due to the QGD, enabling a finite-momentum pairing state.
As we will demonstrate in the following sections, by examining a flat-band model with tunable spin-dependent quantum metrics, we find that even in the flat-band limit, the finite momentum $\bm{Q}$ can be stabilized when pairing is frustrated by QGD, specifically when $\mathcal{G}_{\uparrow}\neq\mathcal{G}_{\downarrow}$ [see Fig. 1(b)], where $\mathcal{G}_{\uparrow}$ and $\mathcal{G}_{\downarrow}$ denote the quantum metrics of the Bloch functions of the paired electrons. Our findings provide a viable route toward realizing flat-band FFLO states.
Effective action from band projection.—To investigate the FFLO instability on the flat band, we first derive the static superconducting Gaussian fluctuations using the band-projection formalism Setty et al. (2023); Iskin (2023); Chen and Law (2024), and this analysis is rather general when considering conventional pairing. Without loss of generality, we consider the local Hubbard interaction $H_{\mathrm{int}}=-U\sum_{i\alpha}\hat{c}^{\dagger}_{i\alpha\uparrow}\hat{c}^{\dagger}_{i\alpha\downarrow}\hat{c}_{i\alpha\downarrow}\hat{c}_{i\alpha\uparrow}$, where $U>0$, $\alpha$ and $i$ denote the orbital and site index.
In order to analyze the low-energy behavior of superconducting pairings on the relevant isolated band using field theory, after the Fourier transformation ($\hat{c}_{i\alpha\sigma}=1/\sqrt{N_{c}}\sum_{\bm{k}}e^{i\bm{k}\cdot\bm{r}_{i}}\hat{c}_{\alpha,\bm{k}\sigma}$, where $N_{c}$ is the number of unit cells) and the uniform pairing approximation Törmä et al. (2018); Huhtinen et al. (2022), we project the interaction onto the target flat band by $\hat{c}_{\alpha,\bm{k}\sigma}\rightarrow u_{\bm{k}\sigma}^{*}(\alpha)\hat{a}_{\bm{k}\sigma}$, where $\hat{c}_{\alpha,\bm{k}\sigma}$ and $\hat{a}_{\bm{k}\sigma}$ denotes the orbital and band-basis electron annihilation operators. We can obtain the effective attractive interaction
$H_{\mathrm{int}}=-g\sum_{\bm{q}}\hat{\theta}^{\dagger}_{\bm{q}}\hat{\theta}_{\bm{q}},$ (2)
$\hat{\theta}_{\bm{q}}=$ $\sum_{\bm{k}}\Lambda^{*}(\bm{k},\bm{q})\hat{a}_{-\bm{k}\downarrow}\hat{a}_{\bm{k}+\bm{q}\uparrow},$
in which $\Lambda(\bm{k},\bm{q})=\sum_{\alpha}u_{\bm{k}+\bm{q}\uparrow}(\alpha)u_{-\bm{k}\downarrow}(\alpha)$ is the form factor, which encodes the information of quantum geometry Chen and Law (2024); Hu et al. (2023). Here the effective coupling constant $g\approx U/N_{\mathrm{orb}}N_{c}$, where $N_{\mathrm{orb}}$ is the number of orbitals in each unit cell. For convenience, we naturally introduce the ideal condition such that the attractive interaction strength $g$ is large compared to the bandwidth $W$ of the isolated flat band but small compared to the band gap $\Delta_{\mathrm{gap}}$ ($W\ll g\ll\Delta_{\mathrm{gap}}$). From Eq. (2), we can arrive at an effective one-band description in the functional integral formalism Altland and Simons (2023), and the effective action can be expressed as
$S=\sum_{k,\sigma}\bar{a}_{k\sigma}\left(-i\omega_{n}+\xi_{\bm{k}\sigma}\right)a_{k\sigma}-g\sum_{q}\bar{\theta}_{q}\theta_{q}.$ (3)
Here we have used the notations $k=\left(\bm{k},\omega_{n}\right)$ and $q=\left(\bm{q},\Omega_{m}\right)$ with $\omega_{n}$ $(\Omega_{m})$ being the fermionic (bosonic) Matsubara frequencies. Then we adopt the standard Hubbard-Stratonovich decoupling of Eq. (3) to derive the effective action $S[\Delta_{q}]$ (for details see Supplemental Material Note I sup ) of the superconducting order parameter field $\Delta_{q}$. By retaining the second order terms of $\Delta_{q}$ in $S[\Delta_{q}]$, the Gaussian action around the normal state saddle point $\Delta_{q}=0$ reads
$S_{G}\left[\Delta_{q}\right]=\sum_{q}\Gamma_{q}^{-1}|\Delta_{q}|^{2},$ (4)
where the coefficient $\Gamma_{q}^{-1}=g^{-1}-\chi^{c}_{q}$ is the pairing propagator, and the static pair susceptibility $\chi^{c}_{\bm{q}}\equiv\chi^{c}_{q=(\bm{q},0)}$ reads
$\chi^{c}_{\bm{q}}=\frac{1}{N_{c}}\sum_{\bm{k}}\frac{1-n_{\mathrm{F}}\left(\xi_{\bm{k+q}}\right)-n_{\mathrm{F}}\left(\xi_{\bm{k}}\right)}{\xi_{\bm{k+q}}+\xi_{\bm{k}}}|\Lambda(\bm{k},\bm{q})|^{2}.$ (5)
Here we use the notation as $\xi_{\bm{k}}=\xi_{\bm{k}\uparrow}=\xi_{-\bm{k}\downarrow}$ and $n_{\mathrm{F}}$ is the Fermi-Dirac distribution. The number equation near the critical temperature is $N=-\beta^{-1}\partial S_{G}[\Delta_{q}=0]/\partial\mu$, where $N$ is the total particle number. This yields $2\mathcal{A}\nu=\sum_{\bm{k}}\left[1-\tanh\left(\beta\xi_{\bm{k}}/2\right)\right]$, where $\mathcal{A}$ is the area of the first Brillouin zone, $\beta^{-1}=k_{B}T$, and $\nu$ is the filling factor of the target band ($0\leq\nu\leq 1$).
Critical temperature of FFLO instability.—By solving the linearized gap equation $g\chi^{c}_{\bm{q}}=1$ together with the number equation, we can derive the critical temperature $T_{c\bm{q}}$ for the instability of a static, spatially non-uniform pairing state characterized by a finite $\bm{q}$ index. Restricting the analysis to single-$\bm{q}$ states, $\Delta_{\bm{q}}\propto\delta_{\bm{q},\bm{Q}}$, the value of $\bm{Q}$ is found at the highest $T_{c\bm{q}}$. If $\bm{Q}\neq 0$ ($\bm{Q}=0$), the pairing instability is associated with the FFLO (BCS) phase.
Importantly, in the flat-band limit, we find $\chi^{c}_{\bm{q}}=\beta\sum_{\bm{k}}|\Lambda(\bm{k},\bm{q})|^{2}(1/2-\nu)/[N_{c}\ln(\nu^{-1}-1)]$. Rewriting Eq. (1) as $\bar{d}^{2}_{\bm{k}+\bm{q},\bm{k}}=1-|\Lambda(\bm{k},\bm{q})|^{2}$, $T_{c\bm{q}}$ with fixed band filling $\nu$ and coupling constant $g$ can be derived as
$T_{c\bm{q}}/T_{0}=1-\bar{d}^{2}_{\mathcal{A}}(\bm{q}),$ (6)
in which the average of the AQD over the Brillouin zone within the band is defined as $\bar{d}^{2}_{\mathcal{A}}(\bm{q})=\sum_{\bm{k}}\bar{d}^{2}_{\bm{k}+\bm{q},\bm{k}}/N_{c}$, and $T_{0}$ is a constant which is independent of $\bm{q}$. The highest critical temperature should be determined by $T_{c\bm{Q}}=T_{0}\{1-\mathrm{min}[\bar{d}^{2}_{\mathcal{A}}(\bm{q})]\}$ over different $\bm{q}$ in the first Brillouin zone, and the corresponding optimal momentum is $\bm{q}=\bm{Q}$. We have clearly seen that the flat-band FFLO instability associated with finite-momentum $\bm{Q}$ is solely determined by the quantum geometric quantity [$\bar{d}^{2}_{\mathcal{A}}(\bm{q})$] from Eq. (6), and is independent of the interaction strength and band filling. This is the main result of our work.
In a $\mathcal{T}$-symmetric system, $u_{\bm{k}\uparrow}=u^{*}_{-\bm{k}\downarrow}$ renders $\bar{d}_{\bm{k}+\bm{q},\bm{k}}$ being the ordinary quantum distance. Since $\bar{d}^{2}_{\bm{k},\bm{k}}=0$, the quantity $\bar{d}^{2}_{\mathcal{A}}(\bm{q})$ always has a global minimum at $\bm{q}=0$. However, as we will discuss, the presence of QGD explicitly breaks the $\mathcal{T}$ symmetry and shifts the large pair susceptibility away from $\bm{q}=0$, allowing the global minimum of $\bar{d}^{2}_{\mathcal{A}}(\bm{q})$ to occur at $\bm{q}\neq 0$. Therefore, with Eq. (6) in hand, it becomes clear that the concept of QGD is both natural and useful for stabilizing the FFLO state in an isolated flat band.
Model Hamiltonian.—To study the Cooper pairings with QGD, now we examine a concrete model Hofmann et al. (2022, 2023); Mao and Chowdhury (2023); Thumin and Bouzerar (2024) that allows for continuous tuning of the quantum geometry. As shown in Fig. 2(a), the $\zeta$-lattice model features two orbitals (A and B) per site on a square lattice, described by the Hamiltonian $H=\sum_{\bm{k}\sigma}\hat{c}_{\bm{k}\sigma}^{\dagger}(\mathcal{H}_{\bm{k}\sigma}-\mu_{\sigma})\hat{c}_{\bm{k}\sigma}$, where
$\mathcal{H}_{\bm{k}\sigma}(\zeta_{\sigma})=-t\left[\lambda_{x}\sin(\zeta_{\sigma}p_{\bm{k}})+m_{\sigma}\lambda_{y}\cos(\zeta_{\sigma}p_{\bm{k}})\right],$ (7)
with $p_{\bm{k}}=\cos k_{x}+\cos k_{y}$. For brevity we set the lattice constant $a=1$ and $\mu_{\uparrow}=\mu_{\downarrow}=\mu$. With the spin index $\sigma=\uparrow,\downarrow$, the fermion operators are defined as $\hat{c}_{\bm{k}\sigma}=(\hat{c}_{A,\bm{k}\sigma},\hat{c}_{B,\bm{k}\sigma})^{T}$. The Pauli matrices $\lambda_{i}$ act on orbital space, and time-reversal flavors are related by $m_{\sigma=\uparrow/\downarrow}=\pm 1$. The periodic function $\zeta_{\sigma}p_{\bm{k}}$ contains spin-dependent long range hoppings [see Fig. 2(a)]. As plotted in Fig. 2(b), this model Hamiltonian possesses a pair of ideal isolated flat bands with band gap $\Delta_{\mathrm{gap}}=2t$. For each spin flavor, the band dispersion $\xi_{\bm{k}\sigma}$ as well as Bloch function $u_{\bm{k}\sigma}$ can be solved as $\xi_{\bm{k}\sigma}=\pm t-\mu$ and
$u_{\bm{k}\sigma}=\frac{1}{\sqrt{2}}\begin{pmatrix}\pm 1\\ e^{im_{\sigma}\zeta_{\sigma}p_{\bm{k}}}\end{pmatrix}.$ (8)
Given the Bloch function $u_{\bm{k}\sigma}$, the spin-dependent quantum metrics can be expressed as $\mathcal{G}_{ab}(\bm{k},\sigma)=\zeta_{\sigma}^{2}\sin(k_{a})\sin(k_{b})/4$. When considering electrons with opposite momentum $\bm{k}$ that yield spin-singlet pairing, the time-reversal partners $(\bm{k},\uparrow)$ and $(-\bm{k},\downarrow)$ form a Cooper pair. Crucially, $(\bm{k},\uparrow)$ and $(-\bm{k},\downarrow)$ are degenerate in energy as $\xi_{\bm{k}\uparrow}=\xi_{-\bm{k}\downarrow}$, but are polarized in quantum metrics when $\zeta_{\uparrow}\neq\zeta_{\downarrow}$, meaning $\mathcal{G}(\bm{k},\uparrow)\neq\mathcal{G}(-\bm{k},\downarrow)$. In this case, $\mathcal{T}$ symmetry of the system is explicitly broken, which is evidenced by the $u_{\bm{k}\uparrow}\neq u_{-\bm{k}\downarrow}^{*}$, rather than by an energy difference. For the $\zeta$-lattice model, we define the dimensionless parameter $\eta=|\zeta_{\uparrow}-\zeta_{\downarrow}|/(\zeta_{\uparrow}+\zeta_{\downarrow})$ as a global measure of QGD. In the following, we will illustrate the role of QGD in the formation of flat-band FFLO states.
Figure 2: (a) The $\zeta$-lattice with spin-dependent long-range hoppings. $\zeta_{\sigma}$ denotes the quantum metric of spin-$\sigma$ sector. (b) The energy dispersions in flat-band limit ($W=0$) and finite bandwidth ($W\neq 0$). (c) $T_{c\bm{Q}}(\eta)$ (dashed blue), $T_{c0}(\eta)$ (solid red), and the corresponding $Q(\eta)$ (solid green) from Eq. (6) as a funtion of $\eta$. The orange dot denotes the Lifshitz point. The $\bm{Q}(\eta)$ calculated from the self-consistent mean-field (MF) theory is plotted in the dashed green line as a comparison. (d) Calculated $T_{c\bm{q}}(\eta)/T_{0}$ from Eq. (6) on $\bm{q}=q(\hat{x}+\hat{y}),\,q\in\left[0,\pi\right]$, for $\eta=0$ and $\eta=0.25$, respectively. In our calculations, we set $\zeta_{\downarrow}=3$ and $\zeta_{\uparrow}=\zeta_{\downarrow}(1+\eta)/(1-\eta)$.
QGD-stabilized flat-band FFLO state.—In terms of Eq. (6), in Fig. 2(c) we plot $T_{c\bm{Q}}(\eta)$ (dashed blue) as well as the corresponding $\bm{Q}(\eta)$ (solid green). As a comparison, $T_{c0}(\eta)$ (solid red) is also depicted. In the BCS phase where $\eta<\eta_{c}=0.22$, $T_{c\bm{Q}}(\eta)$ is identical to $T_{c0}(\eta)$, and $\bm{Q}(\eta)=0$. However, when $\eta>\eta_{c}$, $T_{c\bm{Q}}(\eta)\neq T_{c0}(\eta)$, and a sharp transition occurs in $\bm{Q}(\eta)$ [from $Q(\eta_{c}-)=0$ to $Q(\eta_{c}+)\approx 0.31\pi$], implying a transition from BCS to FFLO phase. The tri-critical point (orange dot) of BCS, FFLO, and the normal phase is the so-called Lifshitz point, which is located at $(\eta_{c}=0.22,T=0.57T_{0})$. We also find that $\bm{Q}$ is four-fold degenerate, which are located at ±$\pm Q(\hat{x}+\hat{y})$ and ±$\pm Q(-\hat{x}+\hat{y})$ sup , so in our calculations we take $\bm{Q}=Q(\hat{x}+\hat{y})$ throughout this work.
To further understand the QGD-driven FFLO transition and the role of AQD, we plot the $T_{c\bm{q}}(\eta)$ evaluated from Eq. (6) in Fig. 2(d), for $\eta=0$ (pink curve) and $\eta=0.25$ (orange curve), respectively. The highest critical temperature is determined by the minimum of $\bar{d}^{2}_{\mathcal{A}}(\bm{q})$, which corresponds to the maximum of $T_{c\bm{q}}$. As we mentioned before, $\mathcal{T}$ symmetry ($\eta=0$) ensures the favorable BCS pairing state ($Q=0$) because of the global maximum of $T_{c\bm{q}}$ always occurs at $q=0$ (dashed green circle on the pink curve). Note that a local maximum of $T_{c\bm{q}}(0)$ can be seen at nonzero $q$ in Fig. 2(d). As QGD ($\eta\neq 0$) decreases $T_{c0}(\eta)$ significantly (the green dashed circle on the orange curve), it is possible that a nonzero $Q$ can be stabilized. Upon increasing $\eta$, the local maximum finally transitions to global maxima (the cyan dashed circle on the orange curve) at $Q(\eta)\neq 0$, where a FFLO phase can naturally manifest. Here our calculations are based on the square lattice model, and in Supplemental Material Note III sup we also show a triangular lattice model, where our theory still applies.
Self-consistent mean field calculations—The pair susceptibility analysis enables us to find FFLO instabilities in the single-band description. Here we numerically show, based on self-consistent mean field study Kitamura et al. (2022), there is a first-order phase transition from BCS to FFLO phase driven by QGD within the previously introduced $\zeta$-lattice model [Eq. (S23)]. Near the critical temperature, the numerical results perfectly match Eq. (6). The mean-field decoupling of the attractive Hubbard interaction yields orbital-dependent order parameters given by $\Delta_{i\alpha}=-U\langle\hat{c}_{i\alpha\downarrow}\hat{c}_{i\alpha\uparrow}\rangle=\Delta_{\alpha}(\bm{q})e^{i\bm{q}\cdot\bm{r}_{i}}$ with $\bm{q}$ index. In the canonical ensemble, the Helmholtz free energy density $F$ is written as
$F(\bm{q})=$ $-\frac{1}{\beta N_{c}}\sum_{\bm{k}}\operatorname{Tr}\left[\operatorname{ln}\left(1+e^{-\beta H_{\mathrm{BdG}}(\bm{k},\bm{q})}\right)\right]$ (9)
$+\frac{1}{U}\sum_{\alpha}|\Delta_{\alpha}(\bm{q})|^{2},$
where the Bogoliubov–de Gennes (BdG) Hamiltonian $H_{\mathrm{BdG}}(\bm{k},\bm{q})$ reads
$H_{\mathrm{BdG}}(\bm{k},\bm{q})=\left(\begin{array}[]{cc}\mathcal{H}_{\bm{k}+\bm{q}\uparrow}-\mu_{\bm{q}}&\hat{\Delta}_{\bm{q}}\\ \hat{\Delta}^{\dagger}_{\bm{q}}&-\mathcal{H}_{-\bm{k}\downarrow}^{*}+\mu_{\bm{q}}\end{array}\right),$ (10)
with the order parameters $\hat{\Delta}_{\bm{q}}=\mathrm{diag}[\Delta_{A}(\bm{q}),\Delta_{B}(\bm{q})]$. $\hat{\Delta}_{\bm{q}}$ and $\mu_{\bm{q}}$ can be determined self-consistently by solving the gap and number equations, respectively. Substituting $\hat{\Delta}_{\bm{q}}$ and $\mu_{\bm{q}}$ into Eq. (9), we can obtain $F(\bm{q})$ for each $\bm{q}$ (see more details in Supplemental Material Note II sup ).
Figure 3: The BCS-FFLO phase diagram, without (a) and with (b) band dispersion. The dashed black line labels $T_{c0}(\eta)$, while the solid red line labels $T_{c\bm{Q}}(\eta)$. The BCS and FFLO phases are colored in light blue and yellow, respectively. Parameters used here: $U=0.4t$, $\nu=0.5$, $\zeta_{\downarrow}=3$, $W/U=0$ in (a) and $W/U=0.12$ in (b). (c) (d) (e): The free energy density landscapes of three highlighted dots in (a) at $T=0.1T_{0}$ and different $\eta$ ($\eta=0.10$, $0.23$, and $0.29$, respectively).
The results from mean-field calculations are summarized in Fig. 3. As shown in Fig. 3(a), the BCS as well as FFLO phases are colored in light blue and yellow, respectively. $T_{c0}(\eta)$ and $T_{c\bm{Q}}(\eta)$ are plotted in the dashed black and the solid red curves. The most favored pairing momentum $\bm{Q}(\eta)$ is independent of the temperature, resulting in a straight line for the boundary between the BCS and FFLO phases, which implies that the phase transition can only be driven by QGD, rather than the thermal fluctuations. The calculated $T_{c0}(\eta)$, $T_{c\bm{Q}}(\eta)$ and $\bm{Q}(\eta)$ [$\bm{Q}(\eta)$ is also plotted in Fig. 2(a) to compare with the one evaluated from the band projection method] exhibits quantitatively good agreements with the corresponding analytical results in Fig. 2(c). To provide a clear picture of the phase transition, in Fig. 3(c) to (e), we show the landscapes of free energy $F(\bm{q})$ for the three points (blue, green, and pink dots) in Fig. 3(a). Initially, in the BCS region ($\eta<\eta_{c}$) [Fig. 3(c)], the global minimum of the free energy $F(\bm{q})$ is at $\bm{q}=0$, which is stable under weak fluctuations. Some local minima exist at $\bm{q}\neq 0$ but are metastable. When $\eta>\eta_{c}$ [Fig. 3(d)], $\bm{q}=0$ turns into a local minimum but remains a metastable state, while some $\bm{q}=\bm{Q}$ points become global minima (the model respects $C_{4}$ rotational symmetry, so the minima have four-fold degeneracy in the whole Brillouin zone, see Supplemental Material Note III sup ), the FFLO phase can manifest. This indicates that the QGD energetically penalizes BCS pairing more than FFLO pairing. A first-order phase transition from BCS to FFLO occurs near $\eta_{c}$, which is consistent with the results obtained from the band-projection method. As $\eta$ continues to increase [Fig. 3(e)], the local minimum at $\bm{q}=0$ eventually nearly disappears.
Robustness at finite bandwidth.—The QGD-induced BCS-FFLO phase transition has been examined in the ideal flat-band case, both analytically and numerically. However, a natural concern arises about whether the theory will break down at finite band dispersion. To show the robustness of QGD-induced BCS-FFLO transition with finite band dispersion, we modify the Bloch Hamiltonian in Eq. (S23) as $\mathcal{H}^{\prime}_{\bm{k}\sigma}=\mathcal{H}_{\bm{k}\sigma}-W\left(\cos k_{x}+\cos k_{y}\right)\lambda_{0}/4$, where an additional nearest hopping term brings a finite bandwidth $W$ as shown in Fig. 2(b). In this case, the Bloch wave functions remain unchanged. We keep $W\ll U$ to ensure a narrow bandwidth compared to the interaction. In Fig. 3(b), the calculated phase diagram at $W/U=0.12$ ($U=0.4t$) is shown. Obviously, compared to the exactly flat band case, the transition boundary line is temperature-dependent and becomes curved due to the nonzero $W$, requiring a larger QGD to reach the FFLO region compared with Fig. 3(a), especially at lower temperatures. This fact indicates that increasing the temperature may also drive the BCS-FFLO phase transition (e.g., along the dashed pink line at $\eta=0.25$ in Fig. 3(b)). The reason is, in Eq. (5), the form factor gets Fermi-surface averaged for a dispersive band, which weakens the effects of QGD. And the thermal excitation energy $k_{B}T$ influences the Fermi-surface average in Eq. (5), reshaping the transition boundary line. On the other hand, for conventional superconductors with a hkighly dispersive band, we have $W\gg U\gg k_{B}T$ in the weak coupling regime, where the QGD-driven mechanism for the FFLO state may not be effective (more discussions, see Supplemental Material Note II sup ), which is consistent with our intuition.
Conclusion and Discussion.—In this work, we have identified the QGD as the quantum-geometric origin of the FFLO pairing phase for flat-band electrons. This mechanism differs from other pair-breaking mechanisms, such as the Zeeman effects in helical superconductors Daido et al. (2022) or the orbital effects in layered superconductors Liu (2017). Although we only consider the spin-singlet pairing channel, there is no obstacle in generalizing it to other pseudo-spin degrees of freedom (such as valley and layer) and unconventional pairing channels.
Besides, throughout this work we expect the LO-type pairing state to be more stable than the FF type due to the ±$\pm\bm{Q}$ degeneracy (see Fig. 3) for the inversion-symmetric model Hamiltonian [Eq. (S23)]. By turning on an extra inversion-breaking term, the ±$\pm\bm{Q}$ degeneracy is lifted, resulting in the FF-type pairing state and emergent superconducting diode effect (for more details, see Supplemental Material IVsup ), driven by QGD. This is an analog to the helical superconductors with Rashba spin-orbit interaction and in-plane magnetic fields Smidman et al. (2017); Kinnunen et al. (2018).
Several possible routes can manifest QGD in solid-state systems. For example, optical cavities have been proposed to engineer the topology and quantum geometry of the electronic states Hübener et al. (2021); Schlawin et al. (2022). Another option is that in ultracold atom systems, the hopping parameters can be controlled artificially Jotzu et al. (2015); Kinnunen et al. (2018). These schemes remain the topics for future study.
Acknowledgements—We thank Patrick A. Lee, Xiao-Gang Wen, Xilin Feng, Ying-Ming Xie, Akito Daido, and Qing-Dong Jiang for inspiring discussions. K. T. L. acknowledges the support of the Ministry of Science and Technology, China, and Hong Kong Research Grant Council through Grants No. 2020YFA0309600, No. RFS2021-6S03, No. C6025-19G, No. AoE/P-701/20, No. 16310520, No. 16307622, and No. 16309223.
Supplemental Material for “Flat-band FFLO State from Quantum Geometric Discrepancy”
Zi-Ting Sun,1 Ruo-Peng Yu,1 Shuai A. Chen,1 Jin-Xin Hu,1 K. T. Law1
1Department of Physics, Hong Kong University of Science and Technology, Clear Water Bay, Hong Kong, China
I I. EFFECTIVE ACTION IN THE BAND-PROJECTION FORMALISM
To study the physics in a multi-band system with pairing instability, it is convenient to project the attractive interaction onto the target flat band. In this section, we will derive the effective action for the single-band description and we will show the full numerical results in the next section. To begin with, we consider a general local attractive Hubbard interaction:
$\hat{H}_{\mathrm{int}}=-U\sum_{i\alpha}\hat{c}_{i\alpha\uparrow}^{\dagger}\hat{c}_{i\alpha\downarrow}^{\dagger}\hat{c}_{i\alpha\downarrow}\hat{c}_{i\alpha\uparrow}.$ (S1)
Here $U$ denotes the attractive interaction strength which can give rise to Cooper pairing states. $i$ and $\alpha$ are the site and orbital index, respectively. The fermion operator $c_{\alpha,\bm{k}\sigma}^{\dagger}$ in the momentum space can be expressed by the Fourier transform:
$\hat{c}_{\alpha,\bm{k}\sigma}^{\dagger}$ $=\frac{1}{\sqrt{N_{c}}}\sum_{i}e^{-i\bm{k}\cdot\bm{r}_{i\alpha}}\hat{c}_{i\alpha\sigma}^{\dagger}$ (S2)
$\hat{c}_{i\alpha\sigma}^{\dagger}$ $=\frac{1}{\sqrt{N_{c}}}\sum_{\bm{k}}e^{i\bm{k}\cdot\bm{r}_{i\alpha}}\hat{c}_{\alpha,\bm{k}\sigma}^{\dagger}.$ (S3)
In this way, we can rewrite the interacting Hamiltonian in the momentum space as
$\hat{H}_{\mathrm{int}}=-\frac{U}{N_{c}}\sum_{\alpha\bm{k}\bm{k}^{\prime}\bm{q}}\hat{c}_{\alpha,\bm{k}+\bm{q}\uparrow}^{\dagger}\hat{c}_{\alpha,-\bm{k}\downarrow}^{\dagger}\hat{c}_{\alpha,-\bm{k}^{\prime}\downarrow}\hat{c}_{\alpha,\bm{k}^{\prime}+\bm{q}\uparrow}.$ (S4)
For the convenience of the analytical treatment, we approximate the interaction term Törmä et al. (2018); Huhtinen et al. (2022); Peotta et al. (2023) as
$\hat{H}_{\mathrm{int}}\approx-\frac{U}{N_{orb}N_{c}}\sum_{\alpha\beta\bm{k}\bm{k}^{\prime}\bm{q}}\hat{c}_{\alpha,\bm{k}+\bm{q}\uparrow}^{\dagger}\hat{c}_{\alpha,-\bm{k}\downarrow}^{\dagger}\hat{c}_{\beta,-\bm{k}^{\prime}\downarrow}\hat{c}_{\beta,\bm{k}^{\prime}+\bm{q}\uparrow}.$ (S5)
We use $g=U/(N_{orb}N_{c})$ to characterize the effective interaction. We project the interaction onto the target flat band by $\hat{c}_{\alpha,\bm{k}\sigma}\rightarrow u_{\bm{k}\sigma}^{*}(\alpha)\hat{a}_{\bm{k}\sigma}$, where $\hat{a}_{\bm{k}\sigma}$ denotes the band-basis electron annihilation operator. Then from the effective $\hat{H}_{\mathrm{int}}$, on the target flat band, the effective action after the band projection can be expressed as
$S=\sum_{k,\sigma}\bar{a}_{k\sigma}\left(-i\omega_{n}+\xi_{\bm{k}\sigma}\right)a_{k\sigma}-g\sum_{q}\bar{\theta}_{q}\theta_{q}.$ (S6)
where
$\theta_{q}=\sum_{k}\Lambda^{*}(\bm{k},\bm{q})a_{-k\downarrow}a_{k+q\uparrow},$ (S7)
in which $\Lambda(\bm{k},\bm{q})=\sum_{\alpha}u_{\bm{k}+\bm{q}\uparrow}(\alpha)u_{-\bm{k}\downarrow}(\alpha)=\langle u_{\bm{k}+\bm{q}\uparrow}|u^{*}_{-\bm{k}\downarrow}\rangle$ is introduced as the form factor. Here we have used the notations $k=\left(\bm{k},\omega_{n}\right)$ and $q=\left(\bm{q},\Omega_{m}\right)$ with $\omega_{n}$ $(\Omega_{m})$ being the fermionic (bosonic) Matsubara frequencies.
Figure S1: (a)-(c): The $\Delta-T$ relation for the $\eta=0.1$ (in BCS region) states paired with the finite momenta of $\bm{q}=q(\hat{x}+\hat{y})$, and $q=0$, $0.19\pi$, $0.24\pi$ respectively. The favoured finite-momentum $Q=0$. (d)-(f): The $\Delta-T$ relation for the $\eta=0.23$ (in FFLO region) states paired with the same $q$ in (a)-(c), and the favoured finite-momentum $Q=0.24\pi$. In (a)-(f), $\Delta_{A}$ is donated by the red solid curve and $\Delta_{B}$ is donated by the black dashed curve.
After a standard Hubbard-Stratonovich decoupling, one can find the action in the Nambu basis $\psi_{k,q}=\left(a_{k+q\uparrow},\bar{a}_{-k\downarrow}\right)^{T}$ takes the form as
$S=\sum_{q}\frac{\left|\Delta_{q}\right|^{2}}{g}+\sum_{k,q}\bar{\psi}_{k,q}\hat{\mathbb{G}}^{-1}_{k,q}\psi_{k,q},$ (S8)
where
$\hat{\mathbb{G}}^{-1}_{k,q}=\left(\begin{array}[]{cc}-i\omega_{n}-i\Omega_{m}+\xi_{\bm{k}+\bm{q}\uparrow}&-\Lambda(\bm{k},\bm{q})\Delta_{q}\\ -\Lambda^{*}(\bm{k},\bm{q})\Delta^{*}_{q}&-i\omega_{n}-\xi_{-\bm{k}\downarrow}\end{array}\right).$ (S9)
After integrating out the fermionic field, we can get
$S[\Delta_{q}]=\sum_{q}\frac{\left|\Delta_{q}\right|^{2}}{g}-\operatorname{Tr}\ln\hat{\mathbb{G}}^{-1}_{k,q},$ (S10)
and expanding it up to the second order in $\Delta_{q}$, the Gaussian action around the trivial saddle point $\Delta_{q}=0$ reads
$S_{G}\left[\Delta_{q}\right]=\sum_{q}\Gamma_{q}^{-1}|\Delta_{q}|^{2},$ (S11)
in which the coefficient
$\Gamma_{q}^{-1}=\left.\frac{\delta^{2}S[\Delta_{q}]}{\delta\Delta^{*}_{q}\delta\Delta_{q}}\right|_{\Delta_{q}=\Delta^{*}_{q}=0}=g^{-1}-\chi^{c}_{q}$ (S12)
is the pairing propagator, and the pair susceptibility $\chi^{c}_{\bm{q}}\equiv\chi^{c}_{q=(\bm{q},0)}$ at the static limit ($\Omega_{m}=0$) reads
$\chi^{c}_{\bm{q}}=\frac{1}{N_{c}}\sum_{\bm{k}}\frac{1-n_{\mathrm{F}}\left(\xi_{\bm{k+q}\uparrow}\right)-n_{\mathrm{F}}\left(\xi_{-\bm{k}\downarrow}\right)}{\xi_{\bm{k+q}\uparrow}+\xi_{-\bm{k}\downarrow}}|\Lambda(\bm{k},\bm{q})|^{2}.$ (S13)
Near the critical temperature, we also have the number equations (at the weak coupling limit), which read
$N_{\sigma}=-\beta^{-1}\frac{\partial S[\Delta_{q}=0]}{\partial\mu_{\sigma}}=\frac{1}{2}\sum_{\bm{k}}\left[1-\tanh\left(\frac{\beta\xi_{\bm{k}\sigma}}{2}\right)\right].$ (S14)
In the ideal flat-band limit, we have $\xi_{\bm{k}\sigma}=-\mu$. One can find that $\chi^{c}_{\bm{q}}=\beta\sum_{\bm{k}}|\Lambda(\bm{k},\bm{q})|^{2}(1/2-\nu)/[N_{c}\ln(\nu^{-1}-1)]$, where $\nu$ is the filling factor. At half filling $\nu=1/2$, $\chi^{c}_{\bm{q}}\rightarrow\beta\sum_{\bm{k}}|\Lambda(\bm{k},\bm{q})|^{2}/(4N_{c})$.
Figure S2: The superconducting phase diagram of the $\zeta$-lattice model. The red solid line labels $T_{c0}(\eta)$, the black dashed line labels $T_{c\bm{Q}}(\eta)$, where $\bm{Q}=Q(\hat{x}+\hat{y})$. The BCS regime and the FFLO regime are labeled in light blue and yellow, respectively. Parameters used here: $U=0.4t$, $\zeta_{\downarrow}=3$, $W/U=0.1$ in (a) and $W/U=0.2$ in (b).
II II. SELF-CONSISTENT MEAN-FIELD THEORY
In this section, we provide the details of the mean-field numerical calculations in the main text. To be specific, we do the decoupling of the on-site attractive Hubbard interaction within the mean-field approximation
$\hat{H}_{\mathrm{int}}$ $=U\sum_{i\alpha}\hat{c}_{i\alpha\uparrow}^{\dagger}\hat{c}_{i\alpha\downarrow}^{\dagger}\hat{c}_{i\alpha\downarrow}\hat{c}_{i\alpha\uparrow}$ (S15)
$\approx\sum_{\alpha\bm{k}}\left(\Delta_{\alpha}\hat{c}_{\alpha,\bm{k}+\bm{q}\uparrow}^{\dagger}\hat{c}_{\alpha,-\bm{k}\downarrow}^{\dagger}+h.c.\right)+\frac{N_{c}}{U}\sum_{\alpha}|\Delta_{\alpha}|^{2},$
where we have used $\Delta_{\alpha}=-U\left\langle\hat{c}_{\alpha,-\bm{k}\downarrow}\hat{c}_{\alpha,\bm{k}+\bm{q}\uparrow}\right\rangle$ and $\Delta_{\alpha}=\Delta_{\alpha}(\bm{q})$ can be determined by the self-consistent gap equation for a given $\bm{q}$. The full mean-field Hamiltonian can be rewritten in the matrix form in Nambu basis
^$\hat{\mathcal{H}}_{\mathrm{MF}}=\sum_{\bm{k},\bm{q}}\psi^{\dagger}(\bm{k},\bm{q})H_{\mathrm{BdG}}(\bm{k},\bm{q})\psi(\bm{k},\bm{q})+\sum_{\bm{k},\bm{q}}\operatorname{Tr}[H_{-\bm{k}\downarrow}-\mu(\bm{q})\mathbb{I}_{2\times 2}]+\frac{N_{c}}{U}\sum_{\alpha}|\Delta_{\alpha}|^{2},$ (S16)
in which
$H_{\mathrm{BdG}}(\bm{k},\bm{q})=\left(\begin{array}[]{cc}H_{\bm{k}+\bm{q}\uparrow}-\mu(\bm{q})\mathbb{I}_{2\times 2}&\operatorname{diag}(\Delta_{\alpha})\\ \operatorname{diag}(\Delta_{\alpha}^{\dagger})&-H_{-\bm{k}\downarrow}^{T}+\mu(\bm{q})\mathbb{I}_{2\times 2}\end{array}\right),$ (S17)
where $\operatorname{diag}(\Delta_{\alpha})=\operatorname{diag}(\Delta_{A},\Delta_{B})$, the momentum $\bm{q}=q_{x}\hat{x}+q_{y}\hat{y}$, and the Nambu spinor reads
$\psi(\bm{k},\bm{q})=\left(\hat{c}_{A,\bm{k}+\bm{q}\uparrow},\hat{c}_{B,\bm{k}+\bm{q}\uparrow},\hat{c}_{A,-\bm{k}\downarrow}^{\dagger},\hat{c}_{B,-\bm{k}\downarrow}^{\dagger}\right)^{T}.$ (S18)
The grand potential is defined as
$\Omega[\mu(\bm{q}),\Delta_{\alpha}(\bm{q}),\bm{q}]$ $=-\frac{1}{\beta}\operatorname{log}\mathcal{Z}=-\frac{1}{\beta}\operatorname{ln}\operatorname{Tr}[e^{-\beta\hat{\mathcal{H}}_{\mathrm{MF}}}]$ (S19)
$=\sum_{\bm{k}}\operatorname{Tr}(H_{-\bm{k}\downarrow}-\mu(\bm{q})\mathbb{I}_{2\times 2})+\frac{N_{c}}{U}\sum_{\alpha}|\Delta_{\alpha}|^{2}-\frac{1}{\beta}\sum_{\bm{k}}\operatorname{Tr}\left[\operatorname{ln}\left(1+e^{-\beta H_{\mathrm{BdG}}(\bm{k},\bm{q})}\right)\right]$
$=\sum_{\bm{k}}\operatorname{Tr}(H_{-\bm{k}\downarrow}-\mu(\bm{q})\mathbb{I}_{2\times 2})+\frac{N_{c}}{U}\sum_{\alpha}|\Delta_{\alpha}|^{2}-\frac{1}{\beta}\sum_{n\bm{k}}\operatorname{ln}\left(1+e^{-\beta E_{n}(\bm{k},\bm{q})}\right),$
where $E_{n}(\bm{k},\bm{q})$ is the n-th eigenvalue of the BdG Hamiltonian $H_{\mathrm{BdG}}(\bm{k},\bm{q})$. By minimizing the grand potential with respect to the pairing potential $\Delta_{\alpha}$ we can get the self-consistent equation:
$\frac{\partial\Omega}{\partial\Delta_{\alpha}^{*}}=\frac{N_{c}}{U}\Delta_{\alpha}(\bm{q})+\sum_{\bm{k}}\langle\hat{c}_{\alpha,-\bm{k}\downarrow}\hat{c}_{\alpha,\bm{k}+\bm{q}\uparrow}\rangle=0,$ (S20)
$\Delta_{\alpha}(\bm{q})=-\frac{U}{N_{c}}\sum_{\bm{k}}\langle\hat{c}_{\alpha,-\bm{k}\downarrow}\hat{c}_{\alpha,\bm{k}+\bm{q}\uparrow}\rangle=-\frac{U}{N_{c}}\sum_{\bm{k}}\operatorname{Tr}\left[\mathcal{U}_{\bm{k},\bm{q}}^{\dagger}\mathcal{M}_{\alpha}\mathcal{U}_{\bm{k},\bm{q}}\hat{\rho}\right].$
And the filling factor $\nu$ can be obtained from the thermodynamic relation:
$N=2N_{c}\nu=-\frac{\partial\Omega}{\partial\mu}=\sum_{\bm{k}}\langle\hat{c}_{a,\bm{k}\uparrow}^{\dagger}\hat{c}_{a,\bm{k}\uparrow}+\hat{c}_{a,\bm{k}\downarrow}^{\dagger}\hat{c}_{a,\bm{k}\downarrow}+\hat{c}_{b,\bm{k}\uparrow}^{\dagger}\hat{c}_{b,\bm{k}\uparrow}+\hat{c}_{b,\bm{k}\downarrow}^{\dagger}\hat{c}_{b,\bm{k}\downarrow}\rangle,$ (S21)
$2\nu=1+\sum_{\bm{k}}\operatorname{Tr}\left[\mathcal{U}_{\bm{k},\bm{q}}^{\dagger}\tau_{z}\otimes\mathbb{I}_{2\times 2}\mathcal{U}_{\bm{k},\bm{q}}\hat{\rho}\right],$
where $\mathcal{U}_{\bm{k},\bm{q}}$ is the unitary operator that diagonalizes the BdG Hamiltonian as $\mathcal{U}_{\bm{k},\bm{q}}^{\dagger}H_{\mathrm{BdG}}(\bm{k},\bm{q})\mathcal{U}_{\bm{k},\bm{q}}=\operatorname{diag}(E_{n})$, $\mathcal{M}_{A}=\left(\begin{array}[]{cccc}0&0&1&0\\ 0&0&0&0\\ 1&0&0&0\\ 0&0&0&0\\ \end{array}\right)$, $\mathcal{M}_{B}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&1&0&0\\ \end{array}\right)$, and the density matrix $\hat{\rho}=\frac{1}{2}\operatorname{diag}\left(\frac{1}{1+e^{\beta E_{n}(\bm{k},\bm{q})}}\right)$. In turn, if the particle number (filling factor $\nu$) is fixed, the chemical potential $\mu(\bm{q})$ can be determined from this equation. Finally, the free energy density (up to a constant given by the summation $\frac{1}{Nc}\sum_{\bm{k}}\operatorname{Tr}(H_{-\bm{k}\downarrow})$) is:
$F(\bm{q})=\frac{1}{N_{c}}\left(\Omega[\mu(\bm{q}),\Delta_{\alpha}(\bm{q}),\bm{q}]+\mu(\bm{q})N\right)=\frac{1}{U}\sum_{\alpha}|\Delta_{\alpha}|^{2}-\frac{1}{\beta N_{c}}\sum_{n\bm{k}}\operatorname{ln}\left(1+e^{-\beta E_{n}(\bm{k},\bm{q})}\right).$ (S22)
In our numerical calculations, we consider the trivial irreducible representation as the intra-orbital pairing channel, i.e., $\Delta_{\alpha\beta}\propto\delta_{{\alpha\beta}}$ is picked up as the only relevant pairing channel, with the other pairing channels neglected. We note that in the previous section, the band-projected interaction Hamiltonian is approximated under the uniform pairing approximation. To show this approximation is accurate near the critical temperature in this model, we plot the $\Delta-T$ relation at different QGD $\eta$ and finite-momentum $\bm{q}$ in Fig. S1. For different $\eta$ and $q$, at low temperatures, $\Delta_{A}$ can differ from $\Delta_{B}$. But near the critical temperature, we find $\Delta_{A}=\Delta_{B}$, which satisfies the uniform pairing condition. This fact explains why our critical temperature analysis based on the uniform pairing condition matches the numerical calculations well.
We also provide more numerical results for the case of finite band dispersion. We keep $W\ll U\ll 2t$ to ensure a narrow bandwidth compared to the band gap and the interaction. In Fig. S2 We show the phase diagram at $W/U=0.1$ and $W/U=0.2$. Compared to Fig.3(b) in the main text, it is clear that the area of the FFLO phase is shrinking with the increase of $W/U$. Especially in Fig. S2(b), the FFLO states can only be found in a very narrow region. It is straightforward to infer that if we keep increasing $W/U$, the FFLO states may vanish, which supports our claim in the main text.
Figure S3: Averaged anomalous quantum distance $\bar{d}^{2}_{\eta}(\bm{q})$ for the square and triangular $\zeta$-lattice model, in the absence (presence) of QGD. (a) Square lattice for $\eta=0$. (b) Square lattice for $\eta=0.6$. (c) Triangular lattice for $\eta=0$. (d) Triangular lattice for $\eta=0.6$. Other parameters used here: $\zeta_{\uparrow}=1$, $t=1$.
Figure S4: (a) The superconducting phase diagram of $\zeta$-lattice model with an inversion breaking term $\gamma$. The $\bm{Q}>0$ (weak helical superconducting state) and $\bm{Q}<0$ (strong helical superconducting state) regimes are colored in blue and yellow, respectively. Parameters used here: $U=0.4t$, $\zeta_{\uparrow}=3$, $\gamma=3$. (b) The red curve shows the finite-momentum $\bm{Q}=Q(1,1)$ versus QGD ($\eta$). Meanwhile, the blue curve is the corresponding diode efficiency $\Theta$. When increasing $\eta$ to across the phase boundary $\eta=0.26$, the diode efficiency $\Theta$ is enhanced dramatically, and the sign of $\bm{Q}$ is reversed. (c) and (d) The free energy (red curve) and supercurrent (blue curve) of 2 highlighted points in (a) with the temperature $T=0.1T_{0}$ but different $\eta$ [$\eta=0.24$ in (c) and $0.29$ in (d)]. Both of the free energy and supercurrent are normalized to [-1, 1]. The green point is in the $\bm{Q}>0$ phase, and the red point is in the $\bm{Q}<0$ phase.
III III. A FLAT-BAND MODEL ON TRIANGULAR LATTICE
In the main text, we study a $\zeta$-lattice model on a square lattice, which respects $C_{4}$ symmetry. To further show the generality of our approach, in this section, we show similar results for the triangular lattice that has orbitals ($A$ and $B$) per site. Similarly, we can write down the Bloch Hamiltonian as
$\mathcal{H}_{\bm{k}\sigma}=-t\left(\lambda_{x}\sin\zeta_{\sigma}p_{\bm{k}}+m_{\sigma}\lambda_{y}\cos\zeta_{\sigma}p_{\bm{k}}\right)-\mu_{\sigma}\lambda_{0},$ (S23)
$p_{\bm{k}}=\sum_{j=1}^{3}\cos(\bm{k}\cdot\bm{R}_{j}).$
where $\bm{R}_{1}=a(1,0),\bm{R}_{1}=a(-1/2,\sqrt{3}/2),\bm{R}_{3}=a(-1/2,-\sqrt{3}/2)$. To demonstrate the general validity of our theory in the main text, we calculate the averaged anomalous quantum distance $\bar{d}^{2}_{\eta}(\bm{q})$ for this model in the absence (presence) of QGD in Fig. S3(c) and (d). As a comparison, we also make the plots for the square lattice model presented in the main text in Fig. S3(a) and (b). In all of these figures, the dark blue region denotes the global minimum of $\bar{d}^{2}_{\eta}(\bm{q})$. We can see that in both (a) and (c), when $\eta=0$, the global minimum of $\bar{d}^{2}_{\eta}(\bm{q})$ is at $\bm{q}=0$. But for (b) and (d), the global minima are not at $\bm{q}=0$, which implies the FFLO instability, as we explained in the main text. It is worth noting that in Fig. S3(b) the FFLO states are four-fold degenerate, while in Fig. S3(d) the FFLO states are six-fold degenerate, due to the different crystalline symmetries of the Bloch function. Our theory shows universality for various model Hamiltonians.
IV. FLAT-BAND HELICAL PAIRING STATE AND SUPERCONDUCTING DIODE EFFECT
The helical superconductivity can be realized in noncentrosymmetric superconductors with the aid of in-plane magnetic field Daido et al. (2022). To achieve an analog of the helical pairing state in a flat-band system with QGD, we manually turn on an extra inversion breaking term in the $\zeta$-lattice model by modifying (recall the periodic function is $\alpha_{\bm{k},\sigma}=\zeta_{\sigma}p_{\bm{k}}$)
$\alpha_{\bm{k},\sigma}=\zeta_{\sigma}p_{\bm{k}}+m_{\sigma}\gamma(\sin k_{x}+\sin k_{y}).$ (S24)
Here the term with parameter $\gamma$ explicitly breaks the inversion symmetry in the Bloch wave function such that $\alpha_{-\bm{k},\sigma}\neq\alpha_{\bm{k},\sigma}$, while still preserving the time-reversal symmetry as $\alpha_{\bm{k},\uparrow}=\alpha_{-\bm{k},\downarrow}$ when $\eta=0$. However, when $\eta\neq 0$, this term entirely lifts ±$\pm\bm{Q}$ degeneracy in the free energy [see Fig. S4(d)], resulting in FF-type pairing state. Here the Cooper pair momentum is along the $\Gamma-M$ line as $\bm{q}=q(\hat{x},\hat{y})$.
The mean-field results are summarized in Fig. S4. In this case, the finite momentum pairing state is favored for any nonzero $\eta$. Interestingly, $\bm{Q}(\eta)$ shows a sharp change near a critical $\eta_{c}$, manifested as the sign change [Fig. S4 (b)], which suggests a first-order phase transition. In Fig. S4 (a), we color the $\bm{Q}>0$ and $\bm{Q}<0$ regimes in blue and yellow, respectively. This is an analog to the weak and strong helical superconducting states Daido et al. (2022).
Recently, the superconducting diode effect (SDE), in which the magnitudes of the critical supercurrents differ in opposite directions, has been studied in helical superconductors Daido et al. (2022); Nadeem et al. (2023). Inspired by this, here we also calculate the diode efficiency $\Theta=(J_{c+}-|J_{c-}|)/(J_{c+}+|J_{c-}|)$ in Fig. S4 (b), where $J_{c\pm}$ is the upper and lower critical supercurrent in the $(1,1)$-direction. It shows that a dramatic enhancement of $\Theta$ appears near $\eta_{c}$, which is similar to what has been proposed in Ref. Daido et al. (2022).
We provide some details in evaluating the supercurrent. The supercurrent can be evaluated by $J(q)=2[\partial F(q)]/\partial q$, and the minimized free energy can be calculated from Eq. (S22). To give a clearer picture of the phase transition, in Fig. S4(c) and (d), we show the free energy $F(q)$ (red curves) as well as the supercurrent $J(q)$ (blue curves) for the two points (green and red dots) in Fig. S4(a). In the blue region [Fig. S4 (a)], the global minimum of the free energy $F(q)$ is at $Q>0$ [Fig. S4 (c)], but $|\bm{Q}|$ is small, associated with the small diode efficiency. When $\eta>\eta_{c}$ [Fig. S4 (d)], the pairing state with $Q<0$ becomes the new global minimum. In this case, an enhanced diode efficiency can be observed in Fig. S4(b).